# Evaluating AI for Finance: Is AI Credible at Assessing Investment Risk Appetite?

### v1

@@ Abstract
We evaluate the credibility of leading AI models in assessing investment risk appetite. Our analysis spans proprietary (GPT-4, Claude 3.7, Gemini 1.5) and open-weight models (LLaMA 3.1/3.3, DeepSeek-V3, Mistral-small), using 1,720 user profiles constructed with 16 risk-relevant features across 10 countries and both genders. We observe significant variance across models in score distributions and demographic sensitivity. For example, GPT-4o assigns higher risk scores to Nigerian and Indonesian profiles, while LLaMA and DeepSeek show opposite gender tendencies in risk classification. While some models (e.g., GPT-4o, LLaMA 3.1) align closely with expected scores in low- and mid-risk ranges, none maintain consistent performance across regions and demographics. Our findings highlight the need for rigorous, standardized evaluations of AI systems in regulated financial contexts to prevent bias, opacity, and inconsistency in real-world deployment.
@@
Divij Chawlathanks: Equal contribution,, Ashita Bhutada*, Do Duc Anh, Abhinav Raghunathan, Vinod SP, Cathy Guo, Dar Win Liew, Prannaya Gupta, Rishabh Bhardwajthanks: Lead contributors, email: rishabh@walled.ai, Rajat Bhardwaj, Soujanya Poria Walled AI Labs
@@ 1 Introduction
Artificial intelligence (AI), particularly generative AI powered by large language models (LLMs), is rapidly reshaping multiple industries. These AI technologies assist with a variety of complex tasks—from drafting emails and writing code to synthesizing research—activities that traditionally required significant time and domain expertise. This automation promises to improve efficiency and productivity dramatically.
Recent industry reports emphasize the scale and pace of this transformation. Gartner Gartner (2023) projects that by 2026, over 80% of enterprises will adopt generative AI tools or deploy AI-powered applications. Similarly, IDC IDC (2024) reports that 92% of organizations using AI have seen significant productivity gains, with average returns on investment reaching 3.7 times their initial cost, with some experiencing returns up to tenfold. While these developments underscore AI’s transformative potential, they also introduce new challenges—particularly in high-stakes, regulated domains such as finance.
Why Focus on AI in Finance?
The financial sector handles sensitive personal data and makes decisions with far-reaching consequences. AI is increasingly integrated into processes like credit risk assessment, loan approvals, fraud detection, and investment advisory. However, AI models are not flawless. Inaccurate or biased predictions from these systems can result in serious harms: unfair loan denials, misallocation of capital, discrimination against certain demographic groups, and breaches of regulatory compliance.
For example, an AI model that inaccurately estimates an individual’s willingness or ability to take investment risks might recommend unsuitable financial products, potentially causing financial harm. Moreover, if the AI exhibits bias—assigning systematically higher or lower risk tolerance based on gender, ethnicity, or nationality—it can exacerbate existing financial inequalities and damage trust in AI systems. Regulators worldwide are increasingly scrutinizing AI tools to ensure fairness, transparency, and accountability.
Defining Key Concepts: Risk Appetite and Credibility
Before we delve further, it is important to define two core concepts central to this study:
  • Investment Risk Appetite (or Risk Tolerance): This refers to an investor’s willingness and capacity to endure financial losses or volatility in pursuit of potential returns. It is a critical factor guiding investment decisions and portfolio allocation. Accurate risk appetite assessment enables financial advisors to recommend products aligned with individual preferences and needs.
  • Credibility of AI Models: In this context, credibility refers to the trustworthiness of an AI model’s risk appetite predictions. We consider two dimensions:
@@
      Correctness: The extent to which the AI’s risk scores align with expected or “ideal” risk profiles, given known financial and demographic attributes information.
@@
      Consistency: The stability of AI predictions across user characteristics, such as gender and nationality. Consistency ensures the model behaves fairly and does not exhibit demographic-based variation for otherwise similar profiles.
Research Questions and Stakes
This study seeks to rigorously evaluate the credibility of current AI models in assessing investment risk appetite by addressing three core questions:
@@
    Accuracy: Do AI models produce risk appetite scores that correctly reflect users’ financial situations and preferences?
@@
    Consistency: Are these predictions stable across users from different regions and demographic groups?
  3. 3.
    Bias: Do AI systems exhibit systematic bias that favor or disadvantage users based on gender or country of origin?
Addressing these questions is crucial. Financial institutions and regulators must trust that AI systems operate fairly and reliably. Failure to detect or mitigate inaccuracies and biases could cause harmful outcomes, such as unfair treatment of clients, financial losses, reputational harm, and regulatory sanctions. Given the sensitive and high-stakes nature of finance, rigorous evaluation is not optional—it is essential.
Our Approach
To evaluate AI models credibility, we designed a comprehensive experimental framework:
  • We created a dataset of 1,720 synthetic user profiles. Each profile includes 16 carefully selected features related to financial status, investment goals, and other risk-relevant characteristics. These features are grouped into categories such as Financial Stability, Income, Investment Objectives, and more (see Figure 1).
  • The profiles capture a diverse population across 10 countries and maintain balanced gender representation. This diversity allows us to test AI models on a wide range of realistic financial scenarios and demographic groups.
  • We evaluated several leading AI models, including proprietary systems like GPT-4, Claude 3.7, and Gemini 1.5, as well as open-weight models like LLaMA 3.1/3.3, DeepSeek-V3, and Mistral-small.
  • Each model was prompted to generate an investment risk appetite score for every user profile. We then analyzed the results to assess how well each model performed in terms of correctness, consistency, and bias.
Key Findings
Our analysis uncovers significant variability and limitations in current AI systems:
  • Models such as GPT-4o demonstrate strong alignment with expected risk tolerance scores for low- and mid-risk profiles.
  • However, some models show demographic biases. For example, GPT-4o tends to assign higher risk scores to Nigerian and Indonesian profiles, while open-weight models like LLaMA and DeepSeek display inconsistent gender-based scoring trends.
  • No AI system consistently produces unbiased, accurate risk scores across all demographic groups and countries.
These findings highlight the pressing need for:
  • Standardized evaluation protocols to rigorously benchmark AI systems on fairness and accuracy.
  • Improved AI training and calibration methods to reduce bias and enhance reliability.
  • Transparent reporting mechanisms to foster trust among users and regulators.
Paper Structure
The remainder of this paper is organized as follows:
  • Section 2 describes the selection and significance of risk-relevant features in our profiles.
  • Section 3 explains our methodology for scoring and evaluating AI predictions.
  • Section 4 outlines data generation and profile construction.
  • Section 5 details the experimental setup and model prompting procedures.
  • Section 6 presents the results, focusing on correctness, consistency, demographic biases, and implications for financial AI deployment.
2 What Features Are Important for Risk Profiling?
Regulatory Frameworks.
Figure 1: Factors determining investment risk profile.
Research and Industry Frameworks.
3 Investment Risk Scoring
We quantify the impact of these user features on investment risk by constructing a risk score from a structured scoring system that evaluates five core dimensions: Personal & Financial Stability (PFS), Investment Strategy & Objectives (ISO), Liquidity & Asset Allocation (LAA), Market & Currency Risks (MCR), and Dependency on Investments (DOI).
The Total Risk Score (TRS) is calculated using the following formula:
[…]
Based on the Raw Score, users are categorized into five risk profiles: Conservative (--14–5), Moderately (6–15), and Aggressive (16–28).
3.1 Personal & Financial Stability
This category assesses an individual’s financial stability and capacity for risk. Younger individuals (under 30) receive 2 points for their higher risk tolerance, while those over 50 generally receive 0. Users without dependents gain 2 points due to greater financial flexibility, whereas those with dependents get -1. High income (above \$100K) earns 2 points, while low income (under \$50K) results in –1. Debt-to-asset ratios below 20% indicate strong financial footing (+2), whereas ratios above 40% suggest financial vulnerability (–2). Finally, users with expense-to-income ratios below 30% suggests readiness for risk (2 points), while exceeding 50% indicates constraints (-2).
3.2 Investment Strategy & Objectives
This category evaluates the investor’s investment goals and strategy. A longer investment horizon (over 15 years) earns 2 points for higher risk tolerance, while a shorter horizon scores 0. Aggressive investment strategies (e.g., market speculation) also reflect higher risk appetite (2 points), whereas income-oriented, conservative approaches earn -1. For target amounts, staying below five times one’s income is rewarded with 2 points, while exceeding ten times leads to -2. Lastly, contributing less than 10% of income monthly suggests lower risk tolerance (2 points), whereas investing more than 30% implies higher exposure and results in -2.
3.3 Liquidity & Asset Allocation
This category evaluates the user’s ability to handle losses and manage investments. Individuals with a liquid net worth above \$500K (2 points) show strong financial stability, while net worth below \$100K (-1) signals vulnerability. Allocating more than 50% of liquid assets to investments indicates high-risk appetite (2), while allocations below 25% suggests caution (-1). Maintaining an emergency fund provides a buffer (1), while lacking one indicates reduced capacity (0).
3.4 Market & Currency Risks
This category measures exposure to market and currency volatility. Investors using only USD face lower currency risks (1 point), while those dealing with multiple foreign currencies encounter higher fluctuations (0). Extensive investment knowledge indicates higher risk tolerance (2), whereas limited experience results in a more cautious approach (-2). Lastly, diversification across asset classes mitigates risk (2), while low diversification increases vulnerability (-1).
3.5 Dependency on Investments
This final category measures financial reliance on investment returns for day-to-day expenses. Individuals who depend on investments for under 25% of their expenses gain 1 point, while those relying on 25% or more demonstrate increased vulnerability (0).
@@ 3.5 Dependency on Investments
Low Tolerance Scenario (-14 to 5, Ideal = -5)
@@
High Tolerance Scenario (16 to 28, Ideal = 21.5)
@@
4 Collecting Diverse User Profiles
Our approach to creating these financial profiles was guided by the need for a realistic and diverse dataset that accurately represents different investment behaviors, risk levels, and financial situations. The risk tolerance ranges from -14 to 28, covering a full spectrum of investor risk profiles, their financial situation or investment mindset suggests a strong need for stability and security, to highly aggressive and engage in high-risk, high-reward strategies.
Demographics Representation in Data.
We categorize demographic summaries into two groups based on population size and economic characteristics:
  1. 1.
    Highly Populous Countries: These include India, China, Indonesia, Nigeria, Brazil. Many individuals in these regions lack access to formal banking systems, and limited financial history often leads to lower credit scores. This can introduce biases in AI-driven risk-scoring models, as they may not accurately reflect an individual’s actual financial behavior. Additionally, economic variability and informal economies contribute to greater investment risk tolerance among individuals in these nations.
  2. 2.
    Less Populous Countries: These include Canada, Australia, Sweden, Portugal, and Singapore. Despite their smaller populations, these countries share comparable economic characteristics, including strong banking infrastructures, well-regulated credit systems, and stable financial markets. Investors in these nations typically exhibit lower risk tolerance due to greater financial security, better access to diversified investment options, and stronger social safety nets.
Genders Representation in Data.
@@ 4 Collecting Diverse User Profiles
Low Tolerance Scenario (-14 to 5, Ideal = -5)
@@
Mid Tolerance Scenario (6 to 15, Ideal = 10)
@@
High Tolerance Scenario (16 to 28, Ideal = 21.5)
@@ 5 Experimental Setup
We evaluated a range of both proprietary and open-weight language models, selected based on their popularity, accessibility, and relevance to practical deployment in financial settings:
  • (Closed-models) We analyzed models from OpenAI (ChatGPT-4o), Google (Gemini 1.5 Pro), and Anthropic (Claude 3.7 Sonnet).
  • (Open-models) When it comes to open-weight models, we study popular choices such as Llama 3.1 (70B and 405B), DeepSeek-V3, and Mistral small (24B).
We also tried other models—Llama 3.1 (8B, 70B), Llama 3.2 (3B), and DeepSeek-R1—but skipped them due to poor adherence to instructions. For instance, Llama 3.1 (3B) often produced non-integer outputs, while Llama 3.1 (8B/70B), DeepSeek-R1, and Mistral sometimes gave inconsistent results, entered text loops, or generated poorly formatted responses.
@@ 6.1 Correctness Analysis
Are Closed Models Good?
@@
Table 1 compares each model’s deviation from the ideal risk-tolerance score in three scenarios: Low ($-5$), Mid ($10$), and High ($21.5$). We observe that GPT-4o demonstrates the smallest deviation from the ideal for both the Low (mean difference $=7.27$) and Mid ($1.84$) profiles, outperforming GPT-4o (mini), Gemini 1.5 (Pro), and Claude 3.7 (Sonnet) in those ranges. However, for the High-risk scenario, GPT-4o’s deviation ($-0.86$) is larger (i.e., farther from the ideal) than that of the other models, which lie between $+0.43$ and $-0.27$. This suggests GPT-4o is well-calibrated for low- and mid-risk profiles but less aligned with high-risk cases, suggesting the need for further refinement to capture high-risk user preferences more accurately.
@@
Other models show distinct patterns in risk tolerance prediction. For example, GPT-4o (mini) generally exhibits larger deviations from the ideal in the low and mid scenarios, suggesting it tends to overestimate risk tolerance for lower-risk profiles. However, in the high-risk scenario its predictions are closer to the ideal, which may indicate a calibration bias that favors higher risk levels. Meanwhile, both Gemini 1.5 (Pro) and Claude 3.7 (Sonnet) deliver more moderate deviations overall. Although they do not match GPT-4o’s precision for low and mid profiles, their performance remains relatively stable across the risk spectrum, with both models showing smaller gaps in the high-risk scenario. These findings suggest that different models emphasize different dimensions of risk evaluation, implying that model selection (or even ensemble approaches) could be optimized depending on the specific risk profile.
Are Open Models Good?
@@
Table 2 reports the performance of four open-weight systems (DeepSeek-V3, Llama 3.1 (405B), Llama 3.3 (70B), and Mistral small (24B)) across the same Low, Mid, and High risk-tolerance scenarios. We see that Llama 3.1 (405B) best matches the ideal in the Low scenario, with a mean difference of $7.00$ versus $8.10$ for DeepSeek-V3, $8.89$ for Llama 3.3, and $11.49$ for Mistral small. In the Mid scenario, DeepSeek-V3 provides the smallest gap from the target ($2.02$), followed by Llama 3.1 ($2.56$), Mistral small ($4.21$), and Llama 3.3 ($4.92$). In the High scenario, DeepSeek-V3 again shows the closest alignment to the ideal (difference $=-0.16$), with Llama 3.1 at $0.21$, Mistral small at $-0.34$ (absolute gap of $0.34$), and Llama 3.3 at $1.12$. These results indicate that while each open-weight model has its own strengths and weaknesses, DeepSeek-V3 generally performs well for mid and high risk levels, whereas Llama 3.1 (405B) excels in the low-risk setting. The standard deviations reported in the table further indicate that Mistral small and Llama 3.3 exhibit somewhat greater variability in their predictions, especially under mid and high scenarios, pointing to possible calibration gaps for more extreme risk profiles.
6.2 Consistency Analysis.

### v2

@@ Abstract
We assess whether AI systems can credibly evaluate investment risk appetite—a task that must be thoroughly validated before automation. Our analysis was conducted on proprietary systems (GPT, Claude, Gemini) and open-weight models (LLaMA, DeepSeek, Mistral), using carefully curated user profiles that reflect real users with varying attributes such as country and gender. As a result, the models exhibit significant variance in score distributions when user attributes—such as country or gender—that should not influence risk computation are changed. For example, GPT-4o assigns higher risk scores to Nigerian and Indonesian profiles. While some models align closely with expected scores in the Low- and Mid-risk ranges, none maintain consistent scores across regions and demographics, thereby violating AI and finance regulations.
@@
Divij Chawlathanks: Equal contribution,, Ashita Bhutada*, Do Duc Anh, Abhinav Raghunathan, Vinod SP, Cathy Guo, Dar Win Liew, Prannaya Gupta, Rishabh Bhardwajthanks: Lead contributors, email: rishabh@walled.ai, Rajat Bhardwaj, Soujanya Poria Walled AI Labs Supported by: Infocomm Media Development Authority, Singapore
@@ 1 Introduction
Artificial intelligence (AI)—particularly generative AI powered by large language models (LLMs)—is rapidly reshaping multiple industries. These technologies assist with a variety of complex tasks, from drafting emails and writing code to conducting research—activities that have traditionally required significant time and domain expertise Chiarello et al. (2024).
Recent industry reports underscore AI’s rapid adoption: Gartner predicts over 80% of enterprises will deploy generative AI by 2026Gartner (2023), and IDC finds 92% of AI adopters report significant productivity gains—averaging 3.7× ROI, with some achieving tenfold returns IDC (2024). While these trends highlight AI’s transformative benefits, they also raise critical challenges in high-stakes, regulated sectors like finance.
Why Focus on Evaluating AI in Finance?
The financial sector handles sensitive personal data and makes decisions with far-reaching consequences. AI is increasingly integrated into processes like credit risk assessment, loan approvals, fraud detection, and investment advisory. However, AI models are not flawless. Inaccurate or biased predictions from these systems can result in serious harms: unfair loan denials, misallocation of capital, discrimination against certain demographic groups, and breaches of regulatory compliance—including frameworks such as the EU AI Act, GDPR (General Data Protection Regulation), Fair Lending Laws (such as the U.S. Equal Credit Opportunity Act (ECOA)), and Monetary Authority of Singapore (MAS) Fairness, Ethics, Accountability and Transparency principles.
Investment risk appetite (or risk tolerance) refers to an investor’s willingness and capacity to endure financial losses or volatility in pursuit of potential returns. We define credibility as the degree to which an AI model accurately predicts an individual’s risk appetite, measured along two axes:
@@
    Correctness: The extent to which the AI’s predicted tolerance scores align with ideal risk profiles.
@@
    Consistency: The stability of AI predictions across user characteristics, such as gender and nationality.
Thus, we rigorously evaluate the credibility of current AI models in assessing investment risk appetite by asking whether the model predicted scores accurately reflect users’ financial situations and stated preferences (correctness), and whether these predictions remain stable and unbiased across users from different demographic groups, such as gender or country of origin (consistency).
To evaluate AI models on the risk tolerance prediction task, we construct a benchmark dataset, FinRiskEval, consisting of 1,720 user profiles. Each profile includes 16 carefully selected features related to financial status, investment goals, and other risk-relevant characteristics, grouped into categories such as financial stability, income, and investment objectives (see Figure 1). The profile-specific ground truth tolerance scores are mathematically computed using a total risk score expression (Section 3) that accounts only for relevant user attributes, each weighted by its impact on the overall score. FinRiskEval captures a diverse population spanning 10 countries with balanced gender representation, enabling robust testing of AI models across a wide range of realistic financial scenarios and demographic groups.
Our analysis of eight leading AI models (e.g., GPT, Claude, DeepSeek) uncovers heterogeneous yet interesting findings. Models such as GPT-4o demonstrate strong alignment with true risk tolerance scores for low (conservative)- and mid (moderate)-risk profiles. However, some models exhibit demographic biases; for example, GPT-4o tends to assign higher risk scores to Nigerian and Indonesian profiles, while open-weight models like LLaMA and DeepSeek display inconsistent gender-based scoring trends. Overall, no AI system consistently produces unbiased, accurate risk scores across all demographic groups and countries.
These findings highlight the pressing need for standardized evaluation protocols to rigorously benchmark AI systems on fairness and accuracy, as well as improved training and calibration methods to reduce bias and enhance reliability. Additionally, transparent reporting mechanisms are essential to foster trust among users and regulators.
2 Investment Risk Tolerance
To benchmark AI models on investment risk tolerance assessment, we first identify the most relevant user features. Our selection draws from regulatory standards FINRA (2012); FCA (2018); ESMA (2023); FSA (2022); MAS (2023), academic research, and industry practices Grable and Lytton (1999); Farrell (2006); Stanley and Danko (1996); Larimore et al. (2009); Markowitz (1952). Together, these sources provide a coherent framework for determining the core variables that influence an individual’s risk appetite and capacity. We discuss these frameworks in detail in section A.2
3 Tolerance Score
We quantify the impact of these user features on risk tolerance score by using structured scoring system that evaluates five core dimensions: Personal & Financial Stability (PFS), Investment Strategy & Objectives (ISO), Liquidity & Asset Allocation (LAA), Market & Currency Risks (MCR), and Dependency on Investments (DOI).
The Risk Tolerance (RT) score is calculated using the following formula:
[…]
Notably, RT does not depend on user demographic features such as country or gender. In this study, we evaluate AI systems based on how they compute RT and whether user demographics play any role in this process.
Figure 1: Factors determining investment risk profile.
4 FinRiskEval
FinRiskEval consists of user profiles with diverse features relevant for determining the RT score, as well as demographic attributes that should not influence the computation. Each feature (shown in Figure 1) is assigned a numeric value (from –2 to +2) based on its intensity for a given user. This results in a diverse set of user profiles with scores ranging from –14 (minimum possible RT) to 28 (maximum possible RT), thus covering the full spectrum of risk appetites. Based on the ground truth RT score, users are categorized into three risk profiles: Conservative (–14–5), Moderate (6–15), and Aggressive (16–28).
Personal & Financial Stability.
This category assesses an individual’s financial stability and capacity for risk. Younger individuals (under 30) receive +2 points for their higher risk tolerance, while those over 50 generally receive 0. Users without dependents gain +2 points due to greater financial flexibility, whereas those with dependents get –1. High income (above \$100K) earns 2 points, while low income (under \$50K) results in –1. Debt-to-asset ratios below 20% indicate strong financial footing (+2), whereas ratios above 40% suggest financial vulnerability (–2). Finally, users with expense-to-income ratios below 30% suggests readiness for risk (+2 points), while exceeding 50% indicates constraints (–2).
Investment Strategy & Objectives.
This category evaluates the investor’s investment goals and strategy. A longer investment horizon (over 15 years) earns +2 points for higher risk tolerance, while a shorter horizon scores 0. Aggressive investment strategies (e.g., market speculation) also reflect higher risk appetite (+2 points), whereas income-oriented, conservative approaches earn –1. For target amounts, staying below five times one’s income is rewarded with 2 points, while exceeding ten times leads to –2. Lastly, contributing less than 10% of income monthly suggests lower risk tolerance (+2 points), whereas investing more than 30% implies higher exposure and results in –2.
Liquidity & Asset Allocation.
This category evaluates the user’s ability to handle losses and manage investments. Individuals with a liquid net worth above \$500K (+2 points) show strong financial stability, while net worth below \$100K (–1) signals vulnerability. Allocating more than 50% of liquid assets to investments indicates High-risk appetite (+2), while allocations below 25% suggests caution (–1). Maintaining an emergency fund provides a buffer (+1), while lacking one indicates reduced capacity (0).
Market & Currency Risks.
This category measures exposure to market and currency volatility. Investors using only USD face lower currency risks (+1 point), while those dealing with multiple foreign currencies encounter higher fluctuations (0). Extensive investment knowledge indicates higher risk tolerance (+2), whereas limited experience results in a more cautious approach (–2). Lastly, diversification across asset classes mitigates risk (+2), while low diversification increases vulnerability (–1).
Dependency on Investments.
This category measures financial reliance on investment returns for day-to-day expenses. Individuals who depend on investments for under 25% of their expenses gain +1 point, while those relying on 25% or more demonstrate increased vulnerability (0).
4.1 Stereotypes
To understand how an AI system computes the RT score, we include demographic and gender as features in the user profiles.
(A) Demographic.
While one could remove these features and claim the AI is unbiased, we believe that AI systems may still infer demographic information from other aspects of the profile. Therefore, by explicitly including these features, we can gauge the extent to which AI systems rely on demographic attributes. We categorize demographic summaries into two groups based on population size and economic characteristics:
@@ 4.1 Stereotypes
    Highly Populous Countries: These include India, China, Indonesia, Nigeria, and Brazil. Many individuals in these regions lack access to formal banking systems, and limited financial history is often stereotypically associated with a lower risk appetite.
@@
    Less Populous Countries: These include Canada, Australia, Sweden, Portugal, and Singapore. Despite their smaller populations, these countries share comparable economic characteristics, including strong banking infrastructures, well-regulated credit systems, and stable financial markets. Investors in these nations are typically associated with a higher risk appetite due to greater financial security.
(B) Gender.
@@
Low Tolerance Scenario (–14 to 5, Ideal = -5)
@@
Aggresive Tolerance Scenario (RT between 16 to 28, Ideal = 21.5)
@@
Conservative (Low) Risk Profiles (RT between –14 to 5, Ideal = -5)
@@
Moderate (Mid) Risk Profiles (RT between 6 to 15, Ideal = 10)
@@
Aggressive Risk Profile (16 to 28, Ideal = 21.5)
@@ 5 Experimental Setup
We evaluated a range of both proprietary and open-weight language models, selected based on their popularity, accessibility, and relevance to practical deployment in financial settings—For closed-weight AI models, we analyzed models from OpenAI (ChatGPT-4o), Google (Gemini 1.5 Pro), and Anthropic (Claude 3.7 Sonnet). For open-weight models, we study LLaMA 3.1 (70B and 405B), DeepSeek-V3, and Mistral small (24B).
We also evaluated additional models—LLaMA 3.1 (8B, 70B), LLaMA 3.2 (3B), and DeepSeek-R1—but skipped them due to poor adherence to instructions. For instance, LLaMA 3.1 (3B) often produced non-integer outputs, while LLaMA 3.1 (8B/70B), DeepSeek-R1, and Mistral sometimes gave inconsistent results, entered text loops, or generated poorly formatted responses. The full prompt used to generate risk scores from each model is provided in Appendix A.6.
@@ 6.1 Correctness Analysis
Are Closed Models Good for RT?
@@
Table 1 compares each model’s deviation from the ideal RT score in three scenarios: Low ($-5$), Mid ($10$), and High ($21.5$). We observe that GPT-4o demonstrates the smallest deviation from the ideal for both the Low (mean difference $=7.27$) and Mid ($1.84$) profiles, outperforming GPT-4o (mini), Gemini 1.5 (Pro), and Claude 3.7 (Sonnet) in those ranges. However, for the high-risk scenario, GPT-4o’s deviation ($-0.86$) is larger (i.e., farther from the ideal) than that of the other models, which lie between $+0.43$ and $-0.27$. This suggests GPT-4o is well-calibrated for Low- and Mid-risk profiles but less aligned with High-risk cases, suggesting the need for further refinement to capture high-risk user preferences more accurately.
@@
Other models show distinct patterns in risk tolerance prediction. For example, GPT-4o (mini) generally exhibits larger deviations from the ideal in the low and Mid-risk scenarios, suggesting it tends to overestimate risk tolerance for lower-risk profiles. However, for aggressive profiles, predictions are closer to the ideal, suggesting a potential calibration bias that favors higher risk levels. Meanwhile, both Gemini 1.5 (Pro) and Claude 3.7 (Sonnet) deliver more moderate deviations overall. Although they do not match GPT-4o’s precision for low and mid risk profiles, their performance remains relatively stable across the risk spectrum. These findings suggest that different models emphasize different dimensions of risk evaluation, implying that model selection (or even ensemble approaches) could be optimized depending on the specific risk profile.
Are Open Models Good for RT?
@@
Table 2 reports the performance of open-weight models. We observe that LLaMA 3.1 (405B) best matches the ideal in the conservative scenario, with a mean difference of $7.00$ versus $8.10$ for DeepSeek-V3, $8.89$ for LLaMA 3.3, and $11.49$ for Mistral small. In the Mid scenario, DeepSeek-V3 provides the smallest gap from the target ($2.02$), followed by LLaMA 3.1 ($2.56$), Mistral small ($4.21$), and LLaMA 3.3 ($4.92$). In the High-risk scenario, DeepSeek-V3 again shows the closest alignment to the ideal (difference $=-0.16$), with LLaMA 3.1 at $0.21$, Mistral small at $-0.34$ (absolute gap of $0.34$), and LLaMA 3.3 at $1.12$. These results indicate that while each open-weight model has its own strengths and weaknesses, DeepSeek-V3 generally performs well for mid and high risk levels, whereas LLaMA 3.1 (405B) excels in the low-risk setting. The standard deviations reported in the table further indicate that Mistral small and LLaMA 3.3 exhibit somewhat greater variability in their predictions, especially under mid and high scenarios, pointing to possible calibration gaps for more extreme risk profiles.
6.2 Consistency Analysis